Math 254A. Homework 10 Due Wednesday, 1 November

1. Let N be a positive integer. Find the dimension of the ℚ-vector subspace of ℝ generated by the set \{sin \frac{nπ}{2N} : n ∈ ℤ\}.

2. Exercise 1 on page 106.

Note that, by the book’s conventions, 0 ≤ a_{ν} < p for all ν.

3. Exercise 3 on page 115.

4. Exercise 4 on page 115.